OpenStudy (anonymous):

What is the slope of the line passing through the points (2, 5) and (0, −4)?

pooja195 (pooja195):

\[\LARGE \frac{y_2-y_1}{x_2-x_1}=slope\]

pooja195 (pooja195):

\[\LARGE \frac{-4-5}{0-2}=slope\]

OpenStudy (anonymous):

-9/-2

pooja195 (pooja195):

Correct

OpenStudy (anonymous):

so my answer is 9 over 2

pooja195 (pooja195):

Yes good job :)

OpenStudy (anonymous):

thanks for the medal

pooja195 (pooja195):

yw

OpenStudy (anonymous):

I got one more question

pooja195 (pooja195):

ok